1. For the reaction, N2(g) 3H2(g)2NH3(g). if the rate of formation of ammonia is 0.04 mol/L min, find the rate of disappearance of H2 and N2 during that same time period. 2. For the following reaction, 5Br BrO3 6H3Br2 3H20, the rate of consumption of bromide ion at some time was found to be 0.20 mo/L s. What is the rate of formation of water over the same time period? 3. For the zero-order reaction, 2NH3(g)N2(g) + 3H2(g), k 2.5 x 10 (ar 1 100°C). Determine the reaction rate when, a) NHI。 0.020 M, b) [NH3, mo/L min 0.040 M. P, the rate law is rate [AJ[B]. Complete the 4. For the hypothetical reaction, A + 2B [Al. 0.20 following table (assume each trial is carried out at a different temperature). Blo 0.10 0.20 0.12 Trial k(L/mol-s) rate(mol/L.s) 6.0 x 10-6 5.2 x 10-4 1.2 x 10-5 1.2 x 10- 0.45 5. For the reaction, A P, the following rate data were obtained 0.10 0.02 0.09 0.016 0.08 0.013 0.07 0.01 Rate(mol/L s) Write the (differential) rate law for the reaction.
media%2F7eb%2F7ebcf7d8-f3d5-41cb-94f1-05
media%2Ffc2%2Ffc29dbd9-8a93-4e42-9ce6-05
media%2F358%2F358723fe-8748-4c90-a3e8-66
Please answer as many questions as you can
Thanks
Answer #1

1)

rate of reaction = (1/2)*rate of formation of NH3 = (1/3)*rate of disappearance of H2 = rate of disappearance of N2

USE:

(1/3)*rate of disappearance of H2 = (1/2)*rate of formation of NH3

rate of disappearance of H2 = (3/2)*rate of formation of NH3

rate of disappearance of H2 = (3/2)*0.04 mol/L.min

= 0.06 mol/L.min

Use:

rate of disappearance of N2 = (1/2)*rate of formation of NH3

= (1/2)*0.04 mol/L.min

= 0.02 mol/L.min

2)

rate of reaction = (1/5)*rate of consumption of Bromide = (1/3)*rate of formation of H2O

Use:

(1/3)*rate of formation of H2O = (1/5)*rate of consumption of Bromide

rate of formation of H2O = (3/5)*rate of consumption of Bromide

= (3/5)*0.20 mol/L.s

= 0.12 mol/L.s